What % of acetylsalicylic acid dissociates to make H3O+1 ions in water? Compare aspirin dissociation to acids like KHP and HCl. You will need to do a little bit of research and look into the acid dissociation constant, abbreviated by the symbol Ka. For context and clarity, you will want to define acid/base strength and acid dissociation constant.
Answer #1

HCl is strong acid with high Ka value , Ka = 10^ 6 nearly

KHP has pka of 5.4 which means Ka = 10^ -5.4 = 4 x 10^ -6

Asprin has Ka =  3.16 x 10^ -4

Ka is indication of strength of acid. Hence Asprin is stronger acid than KHP but weaker than HCl.

Now Let initial Molarity of Asprinin is 1M , Let acid denoated by HA

then HA + H2O (l) <---> A- + H3O+

at equilibirum [HA] = 1-X , [A-] =[H3O+] = X

Ka = [H3O+] [A-] / [HA]

3.16 x 10^ -4 = X^2 / ( 1-X) ( since Ka is small ,X is small , hence 1-X is nearly equal to 1)

3.16 x 10^ -4 = X^2 / ( 1)

X = ( 3.16 x 10^ -4) ^ 1/2 = 0.0178 =H3O+]

% of acid dissociated to H3O+ = 100 x [H3O+] / [HA] initial = 100 x ( 0.0178 M) / ( 1M)

= 1.78 %